# Multi-level Temporal-channel Speaker Retrieval for Zero-shot Voice Conversion

### v2

@@ III-C Speech Decoder: Speech Representation Fusion
Figure 5: The cycle-based training process of MTCR-VC. (a) Paired Path. (b) Unpaired Path. Note that speech utterances X and Y come from different speakers and have different linguistic content.
@@ III-D Cycle-based Training Strategy
Since the corresponding ground-truth speech of $Y_{x}$ is unavailable in the unpaired path (Fig 5(b)), only mel reconstruction loss is hard to constrain the training process. We introduce perceptual constraints in three aspects: style, content, and speaker, to supervise the training process and also improve the disentanglement ability. The perceptual constraints are illustrated in Fig. 6. Similarly to Wang et al. [34], we calculate the content loss $\mathcal{L}_{con}$ and style loss $\mathcal{L}_{sty}$ to ensure the consistency of content and style between the converted speech and the source speech. The speaker loss $\mathcal{L}_{spk}$ calculated via the MTCR speaker module measures the global difference of speaker representations between the converted speech and the target speaker’s speech at each layer. The details of the loss functions are introduced in Section III-E.
Figure 6: The perceptual constraints used in the training of MTCR-VC. The block in purple represents pre-trained model, which is fixed without parameter update in the model.
@@ III-E2 Style loss
As shown in Fig. 7, a pre-trained speech emotion recognition(SER) model [49] is used as a style model to calculate the style loss. $\mathcal{L}_{sty}$ is calculated from different levels, which stand for the abstraction degree of the hidden representation to the style. Features $H_{l}$, $H_{m}$, and $H_{h}$ from low- to high-level hidden layers, respectively, are used to obtain the style loss:
@@ III-E3 Content loss
The content loss $\mathcal{L}_{con}$ is generated by a pre-trained content model which has the same architecture as CBHG module [50]. The content model is trained for BNF prediction. During the training of MTCR-VC, the content loss between BNF features $B$ and $\hat{B}$ extracted from the source mel spectrogram and predicted mel spectrogram can be described as:
@@ III-E4 Speaker loss
where $Z_{s_{l}}$ and $\hat{Z_{s_{l}}}$ are obtained from the target speaker’s mel spectrogram and predicted mel spectrogram, respectively. $Avg(\cdot)$ means that average the representation along the temporal axis.
@@ IV-A Datasets
During the training stage, 1,000 speakers from the English dataset LibriTTS [51] are used to train the VC model. In the zero-shot voice conversion stage, we test two types of target speakers, including pre-reserved in-dataset (ID) speakers from LibriTTS and out-of-dataset (OOD) speakers from VCTK [52], CMU Arctic [53], and HiFi-TTS datasets [54]. The OOD testing is particularly used to show the robustness of our approach. Information regarding the recording domain, environment, and signal-to-noise ratio (SNR) of the datasets used in this work can be found in Table I. Additionally, speech utterances from VCTK and CMU Arctic are selected as source speech for the test. We randomly sampled 1,000 testing pairs for ID and OOD speakers. We utilize the open-source conformer ASR model [^2] https://github.com/wenet-e2e/wenet, which is trained on LibriSpeech [55], to extract BNF. The SV model [32] is trained on Voxceleb2 [56]. The style and content models for perceptual constraints are trained on ESD [57] and LibriTTS, respectively. To reconstruct waveform from Mel spectrogram, we utilize a universal vocoder TFGAN [58], which is trained on 1,000 hours of an internal dataset.
@@ IV-B Implementation Details
All speech utterances are re-sampled to 24KHz, and 80-dim Mel spectrogram is computed with 50ms frame length and 10ms frame shift. The ASR encoder output BNF has a dimension of 256. We use Pyworld [^3] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder to extract 1-dim logarithmic-domain fundamental frequency (lf0). The lf0 normalized by utterance-level min-max method is used as the pitch contour input of the pitch encoder. And the down-sampling rate in the pitch encoder is 8. The x-vector from the SV model is a 192-dim feature. In the MTCR speaker module, the scale factors $\gamma_{t}$ and $\gamma_{c}$ are both set to 4, and the temporal ranges $\gamma_{tr}$ from low level (top TCR block in Fig. 2) to high level (bottom TCR block in Fig. 2) are 16, 4, 1, which keep the temporal range at a length of 640ms. During training, we set the loss weights as $\lambda_{mel}=4$, $\lambda_{sty}=0.1$, $\lambda_{con}=0.01$, $\lambda_{spk}=0.1$. We train MTCR-VC for 200 epochs with a batch size of 128. The learning rate is set to 1e-5 and decays every 50,000 steps with a decay rate of 0.5. The prenet in the MTCR speaker module consists of a single convolution layer. The configuration of convolution layers in the MTCR speaker module and fusion blocks of the speech decoder is 3$\times$3 filter with 1$\times$1 stride and replication padding. The speech decoder keeps the same settings for the fusion blocks, smoother, and postnet as [23]. For perceptual constraints, following [34], the style model comprises a reference encoder and 3 FC layers, while the content model adopts a CBHG structure [50].
TABLE II: Results of subjective and objective evaluations for zero-shot VC on ID and OOD speakers.
@@ IV-C Comparison Models
To evaluate the performance of the proposed MTCR-VC on the zero-shot VC task, we compare it with three recent representative zero-shot VC models with different speaker modeling methods, including speaker embedding extracted from the SV, utterance-level speaker representation learning, and multi- level speaker representation learning. Notably, all VC models are trained on the same dataset and the same vocoder is utilized to reconstruct waveform from the Mel spectrogram for all the VC models. Details of these comparison models are introduced as follows.
MediumVC [10] adopts speaker representation extracted from a pre-trained SV model. It divides zero-shot VC (any-to-any) into any-to-one and one-to-any VC by a two-stage strategy. Subsequently, the any-to-one VC first converts the source speech to the speech of a specific non-target speaker, and then one-to-any VC converts the resulting speech to the target speech conditioned on the speaker representation of the target speaker’s speech. We build both the two-stage VC models using a content encoder and a speech decoder, which have similar structures to that of MTCR-VC. However, the cross-attention in the fusion block is removed, and the speaker representation is added in each fusion block.
@@
FragmentVC [23] learns variable-length speaker representations from multiple layers and feeds them to the corresponding decoder layers. We use the officially released open-source code [^5] https://github.com/yistLin/FragmentVC to implement FragmentVC. To ensure a fair comparison, we use BNF as the input instead of wav2vec feature to avoid any potential speaker timbre leakage from the source speech.
@@ IV-D Evaluation Metrics
Objective Metrics. For objective evaluations, we calculate word error rate (WER) via the ASR model to validate the intelligibility of the converted speech. Moreover, we calculate the Pearson correlation coefficient of lf0 between the source speech and converted speech to measure the consistency of speaking style, denoting $\boldsymbol{P_{lf0}}$. As in previous work [23], speaker accuracy (ACC) is calculated using an SV model to measure whether the converted speech is from the target speaker. The accuracy is determined by the cosine similarity between the SV’s embeddings of two speech utterances exceeding a predefined threshold, based on the equal error rate (EER) of the SV model over the considered dataset. Better speaker similarity would result in higher speaker accuracy.
@@ V Experimental Results
This section presents the results of subjective and objective evaluations as well as the ablation study conducted on the proposed model. Additionally, a detailed analysis of the MTCR speaker module is provided.
@@ V-A1 Subjective Evaluation
Table II presents the NMOS and SMOS results for zero-shot VC on ID and OOD speakers. In terms of speech naturalness, our MTCR-VC outperforms SRDVC and FragmentVC, both of which model speaker timbre directly from speech during training. While, Medium-VC gets better naturalness than MTCR-VC, likely due to its use of utterance-level speaker modeling and noise-robust speaker verification (SV) model. Specifically, directly using x-vector extracted from the noise-robust SV model as the speaker representation results in less noisy converted speech and better speech naturalness.
Regarding speaker similarity, SMOS is degraded from ID to OOD speakers for all models. This degradation suggests that the performance is affected due to a mismatch between the training and testing sets. However, MTCR-VC gets better speaker similarity and a smaller SMOS gap (0.02) between ID and OOD speakers, effectively mitigating the mismatch and achieving robust performance in zero-shot VC. Additionally, it can be found that FragmentVC with fine-grained modeling outperforms SRDVC, demonstrating that the coarse utterance-level speaker modeling is insufficient for zero-shot VC. However, without considering the varying speaker information along the temporal and frequency dimensions in the speaker modeling process, FragmentVC is difficult to perform well. On the other hand, MediumVC achieves relatively good speaker similarity than the other two comparison models but suffers from the insufficient representation of x-vector for the perceptual speaker timbre in the sense of human hearing. These findings highlight the superior performance of MTCR-VC for speaker modeling while preserving good speech naturalness.
TABLE III: Results of ablation studies.
@@ V-A2 Objective Evaluation
Table II presents the objective results in the last three columns. As illustrated, similar to the subjective results, the SMOS gap between ID and OOD speakers is also reflected in the objective metric of speaker accuracy. Notably, our proposed MTCR-VC achieves the highest speaker accuracy for both ID and OOD speakers. Additionally, MTCR-VC obtains the best performance on WER and $P_{lf0}$, indicating better speech intelligibility and speaking style consistency. Consistent with subjective results, the objective results also demonstrate that MTCR-VC outperforms other models in speaker similarity while preserving good speech naturalness.
@@ V-B Ablation Study
As shown in Table III, We conduct further studies to assess the effectiveness of MTCR-VC with ablations on the MTCR speaker module, cycle-based training strategy, and perceptual constraints.
@@ V-B1 MTCR Speaker Module
To confirm the efficacy of the MTCR speaker module, we replace it with three convolution layers, as in Lin et al. [23], forming the model w/o MTCR (Conv). Besides, we also implement a model w/o MTCR (SV), which only uses SV embedding to represent speaker timbre, to involve this ablation. As can seen in Table III, discarding the MTCR speaker module leads to a noticeable decrease in speaker accuracy. Furthermore, the absence of the MTCR module weakens the disentanglement ability, resulting in higher WER and lower $P_{lf0}$ scores. These findings indicate the effectiveness of the MTCR speaker module for speaker modeling and its positive impact on zero-shot VC.
@@ V-B3 Perceptual Constraints
Table III shows the effect of perceptual losses we used on MTCR-VC. As can be seen, the constraints of style, content, and speaker play important roles. The dropping of these constraints brings corresponding performance degradation in $P_{lf0}$, WER, and ACC, respectively. The result indicates that the three perceptual constraints effectively improve speech disentanglement performance in the zero-shot VC task. Among them, w/o $L_{spk}$ cause small degradation to speaker accuracy because the cycle-based training process implicitly ensures the consistency of speaker timbre.
@@
TABLE IV: Validation results of multi-level TCR blocks.
@@ V-C1 Speaker Representation
To explore what is learned in the speaker representation, we first visualize the speaker representations extracted from each TCR block with t-SNE. Since the speaker representation is of variable lengths and this visualization can only be done on a single vector, the speaker representation is averaged along the temporal axis to obtain a single vector and then visualized from a global perspective. For this visualization, we select 8 speakers from the testset, with each speaker contributing 200 utterances. As shown in Fig. 8, speaker representations from the low-level TCR block (top TCR block in Fig. 2) group into 8 speaker clusters, indicating the presence of explicit speaker discriminative information. As the layer deepens, the speaker clusters start to overlap, eventually becoming indistinguishable. The low-level block is close to the speech spectrogram and thus can easily capture more speaker timbre information. In contrast, the high-level block captures much finer speaker-related characteristics but less speaker-discriminative information. This phenomenon is consistent with a previous zero-shot study [28]. Please note that this visualization only analyzes the representation from global perspective, but the speaker representations used in the MTCR-VC are time-varying features.
Therefore, to further investigate the impact of different TCR blocks, we objectively evaluate the performance of MTCR-VC by sequentially discarding blocks from the high-level layer to the low-level layer. Specifically, the variant model w/o $Z_{s_{3}}$ only uses the first two TCR blocks in the speaker module and is trained from scratch to verify the role of $Z_{s_{3}}$, compared to the proposed MTCR-VC. Besides, discarding both second and third TCR blocks is also verified, forming the model w/o $Z_{s_{2}}$ $Z_{s_{3}}$. The validation results are presented in Table IV. As can be seen, discarding the use of the third block in the model w/o $Z_{s_{3}}$ leads to decreased speaker accuracy. It suggests that the third block also learns time-varying speaker representation which contains fine-grained speaker information, even though the averaged $Z_{s_{3}}$ shows little discrimination among different speakers from the global perspective in Fig 8. Furthermore, removing the second block and third block in the model w/o $Z_{s_{2}}$ $Z_{s_{3}}$ leads to more performance degradation in speaker accuracy. This indicates that the speaker representations learned in $Z_{s_{2}}$ and $Z_{s_{3}}$ that is important for speaker accuracy in MTCR-VC. Additionally, it can be found that discarding the use of TCR blocks increases the WER. This outcome occurs because the speaker representation participates in the speech representation fusion, and the performance of speaker modeling affects the speech generation process.
Refer to caption (a) Spectrogram of a speaker utterance Refer to caption (b) Pitch contour of the utterance
Refer to caption (c) Channel retrieval Refer to caption (d) Temporal retrieval

### v3

@@ III-C Speech Decoder: Speech Representation Fusion
Figure 5: The cycle-based training process of MTCR-VC. (a) Paired Path. (b) Unpaired Path. Note that speech utterances X and Y come from different speakers with different linguistic content.
@@ III-D Cycle-based Training Strategy
Since the corresponding ground-truth speech of $Y_{x}$ is unavailable in the unpaired path (Fig 5(b)), only mel reconstruction loss is hard to constrain the training process. We introduce perceptual constraints in style, content, and speaker to supervise the training process and improve the disentanglement ability. The perceptual constraints are illustrated in Fig. 6. Similarly to Wang et al. [34], we calculate the content loss $\mathcal{L}_{con}$ and style loss $\mathcal{L}_{sty}$ to ensure the consistency of content and style between the converted speech and the source speech. The speaker loss $\mathcal{L}_{spk}$ calculated via the MTCR speaker module measures the global difference of speaker representations between the converted speech and the target speaker’s speech at each layer. The details of the loss functions are introduced in Section III-E.
Figure 6: The perceptual constraints used in the training of MTCR-VC. The block in purple represents the pre-trained model, which is fixed without parameter update in the model.
@@ III-E2 Style loss
As shown in Fig. 7, a pre-trained speech emotion recognition (SER) model [49] is used as a style model to calculate the style loss. $\mathcal{L}_{sty}$ is calculated from different levels, accounting for the abstraction degree of the hidden representation to the style. Features $H_{l}$, $H_{m}$, and $H_{h}$ from low- to high-level hidden layers, respectively, are used to obtain the style loss:
@@ III-E3 Content loss
The content loss $\mathcal{L}_{con}$ is generated by a pre-trained content model that has the same architecture as the CBHG module [50]. The content model is trained for BNF prediction. During the training of MTCR-VC, the content loss between BNF features $B$ and $\hat{B}$ extracted from the source mel spectrogram and predicted mel spectrogram can be described as:
@@ III-E4 Speaker loss
where $Z_{s_{l}}$ and $\hat{Z_{s_{l}}}$ are obtained from the target speaker’s mel spectrogram and predicted mel spectrogram, respectively. $Avg(\cdot)$ means averaging the representation along the temporal axis.
@@ IV-A Datasets
During the training stage, 1,000 speakers from the English dataset LibriTTS [51] are used to train the VC model. In the zero-shot voice conversion stage, we test two types of target speakers, including pre-reserved in-dataset (ID) speakers from LibriTTS and out-of-dataset (OOD) speakers from VCTK [52], CMU Arctic [53], and HiFi-TTS datasets [54]. The OOD testing is particularly used to show the robustness of our approach. Information regarding the recording domain, environment, and signal-to-noise ratio (SNR) of the datasets used in this work can be found in Table I. Additionally, speech utterances from VCTK and CMU Arctic are selected as source speech for the test. We randomly sampled 1,000 testing pairs for ID and OOD speakers. We use the open-source conformer ASR model [^2] https://github.com/wenet-e2e/wenet, trained on LibriSpeech [55], to extract BNF. The SV model [32] is trained on Voxceleb2 [56]. The style and content models for perceptual constraints are trained on ESD [57] and LibriTTS, respectively. To reconstruct waveform from mel spectrogram, we utilize a universal vocoder TFGAN [58], which is trained on 1,000 hours of an internal dataset.
@@
TABLE II: The differences of the speaker modeling methods used by comparison systems.
Utterance- or Multi-level SV-based Time-varying Channel-aware Learnable
MediumVC U $\times$ $\times$ $\times$
SRDVC U $\times$ $\times$ $\times$
FragmentVC M $\times$ $\times$
MTCR-VC M
@@ IV-B Implementation Details
All speech utterances are re-sampled to 24KHz, and an 80-dim mel spectrogram is computed with a 50ms frame length and 10ms frameshift. The ASR encoder output BNF has a dimension of 256. We use Pyworld [^3] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder to extract 1-dim logarithmic-domain fundamental frequency (lf0). The lf0 normalized by utterance-level min-max method is used as the pitch contour input of the pitch encoder. And the down-sampling rate in the pitch encoder is 8. The x-vector from the SV model is a 192-dim feature. In the MTCR speaker module, the scale factors $\gamma_{t}$ and $\gamma_{c}$ are both set to 4, and the temporal ranges $\gamma_{tr}$ from low level (top TCR block in Fig. 2) to high level (bottom TCR block in Fig. 2) are 16, 4, 1, which keep the temporal range at a length of 640ms. During training, we set the loss weights as $\lambda_{mel}=4$, $\lambda_{sty}=0.1$, $\lambda_{con}=0.01$, $\lambda_{spk}=0.1$, based on empirical experience. We train MTCR-VC for 200 epochs with a batch size of 128. The learning rate is set to 1e-5 and decays every 50,000 steps with a decay rate of 0.5. The prenet in the MTCR speaker module consists of a single convolution layer. The configuration of convolution layers in the MTCR speaker module and fusion blocks of the speech decoder is 3$\times$3 filter with 1$\times$1 stride and replication padding. The speech decoder keeps the same settings for the fusion blocks, smoother, and postnet as [23]. For perceptual constraints, following [34], the style model comprises a reference encoder and 3 FC layers, while the content model adopts a CBHG structure [50].
TABLE III: Results of subjective and objective evaluations for zero-shot VC on ID and OOD speakers.
@@ IV-C Comparison Models
To evaluate the performance of the proposed MTCR-VC on the zero-shot VC task, we compare it with three recent representative zero-shot VC models with different speaker modeling methods, including speaker embedding extracted from the SV, utterance-level speaker representation learning, and multi- level speaker representation learning. The differences of speaker modeling methods used by comparison models are listed in Table II. Notably, all VC models are trained on the same dataset, and the same vocoder is used to reconstruct waveform from the mel spectrogram for all the VC models. Details of these comparison models are introduced as follows.
MediumVC [10] adopts speaker representation extracted from a pre-trained SV model. It divides zero-shot VC (any-to-any) into any-to-one and one-to-any VC by a two-stage strategy. Subsequently, the any-to-one VC first converts the source speech to the speech of a specific non-target speaker, and then one-to-any VC converts the resulting speech to the target speech conditioned on the speaker representation of the target speaker’s speech. We build both the two-stage VC models using a content encoder and a speech decoder, which have structures similar to those of MTCR-VC. However, the cross-attention in the fusion block is removed, and the speaker representation is added in each fusion block.
@@
FragmentVC [23] learns variable-length speaker representations from multiple layers and feeds them to the corresponding decoder layers. We use the officially released open-source code [^5] https://github.com/yistLin/FragmentVC to implement FragmentVC. To ensure a fair comparison, we use BNF as the input instead of the wav2vec feature to avoid any potential speaker timbre leakage from the source speech.
@@ IV-D Evaluation Metrics
Objective Metrics. For objective evaluations, we calculate word error rate (WER) via the ASR model to validate the intelligibility of the converted speech. Moreover, we calculate the Pearson correlation coefficient of lf0 between the source speech and converted speech to measure the consistency of speaking style, denoting $\boldsymbol{P_{lf0}}$. As in previous work [23], speaker accuracy (ACC) is calculated using an SV model to measure whether the converted speech is from the target speaker. The accuracy is determined by the cosine similarity between the SV’s embeddings of two speech utterances exceeding a predefined threshold, based on the SV model’s equal error rate (EER) over the considered dataset. Better speaker similarity would result in higher speaker accuracy.
@@ V Experimental Results
This section presents the results of subjective and objective evaluations and the ablation study conducted on the proposed model. Additionally, a detailed analysis of the MTCR speaker module is provided.
@@ V-A1 Subjective Evaluation
Table III presents the NMOS and SMOS results for zero-shot VC on ID and OOD speakers. Regarding speech naturalness, our MTCR-VC outperforms SRDVC and FragmentVC, both of which model speaker timbre directly from speech during training. Meanwhile, Medium-VC gets better naturalness than MTCR-VC, likely due to its use of utterance-level speaker modeling and noise-robust speaker verification (SV) model. Specifically, directly using the x-vector extracted from the noise-robust SV model as the speaker representation results in less noisy converted speech and better speech naturalness.
Regarding speaker similarity, SMOS is degraded from ID to OOD speakers for all models. This degradation suggests that the performance is affected due to a mismatch between the training and testing sets. However, MTCR-VC gets better speaker similarity and a smaller SMOS gap (0.02) between ID and OOD speakers, effectively mitigating the mismatch and achieving robust performance in zero-shot VC. Additionally, it can be found that FragmentVC with fine-grained modeling outperforms SRDVC, demonstrating that the coarse utterance-level speaker modeling is insufficient for zero-shot VC. However, without considering the varying speaker information along the temporal and frequency dimensions in the speaker modeling process, FragmentVC is difficult to perform well. On the other hand, MediumVC achieves relatively better speaker similarity than the other two comparison models but suffers from the insufficient representation of the x-vector for the perceptual speaker timbre in the sense of human hearing. These findings highlight the superior performance of MTCR-VC for speaker modeling while preserving good speech naturalness.
TABLE IV: Results of ablation studies.
@@ V-A2 Objective Evaluation
Table III presents the objective results in the last three columns. As illustrated, similar to the subjective results, the SMOS gap between ID and OOD speakers is also reflected in the objective metric of speaker accuracy. Notably, our proposed MTCR-VC achieves the highest speaker accuracy for both ID and OOD speakers. Additionally, MTCR-VC obtains the best performance on WER and $P_{lf0}$, indicating better speech intelligibility and speaking style consistency. Consistent with subjective results, the objective results also demonstrate that MTCR-VC outperforms other models in speaker similarity while preserving good speech naturalness.
@@ V-B Ablation Study
As shown in Table IV, We conduct further studies to assess the effectiveness of MTCR-VC with ablations on the MTCR speaker module, cycle-based training strategy, and perceptual constraints.
@@ V-B1 MTCR Speaker Module
To confirm the efficacy of the MTCR speaker module, we replace it with three convolution layers, as in Lin et al. [23], forming the model w/o MTCR (Conv). Besides, we also implement a model w/o MTCR (SV), which only uses SV embedding to represent speaker timbre to involve this ablation. As can be seen in Table IV, discarding the MTCR speaker module leads to a noticeable decrease in speaker accuracy. Furthermore, the absence of the MTCR module weakens the disentanglement ability, resulting in higher WER and lower $P_{lf0}$ scores. These findings indicate the effectiveness of the MTCR speaker module for speaker modeling and its positive impact on zero-shot VC.
@@ V-B3 Perceptual Constraints
Table IV shows the effect of perceptual losses we used on MTCR-VC. As can be seen, the constraints of style, content, and speaker play important roles. Dropping these constraints brings corresponding performance degradation in $P_{lf0}$, WER, and ACC, respectively. The result indicates that the three perceptual constraints effectively improve speech disentanglement performance in the zero-shot VC task. Among them, w/o $L_{spk}$ cause small degradation to speaker accuracy because the cycle-based training process implicitly ensures the consistency of speaker timbre.
@@
TABLE V: Validation results of multi-level TCR blocks.
@@ V-C1 Speaker Representation
To explore what is learned in the speaker representation, we first visualize the speaker representations extracted from each TCR block with t-SNE. Since the speaker representation is of variable lengths and this visualization can only be done on a single vector, the speaker representation is averaged along the temporal axis to obtain a single vector and then visualized from a global perspective. We select 8 speakers from the test set for this visualization, each contributing 200 utterances. As shown in Fig. 8, speaker representations from the low-level TCR block (top TCR block in Fig. 2) group into 8 speaker clusters, indicating the presence of explicit speaker discriminative information. As the layer deepens, the speaker clusters start to overlap, eventually becoming indistinguishable. The low-level block is close to the speech spectrogram and thus can easily capture more speaker timbre information. In contrast, the high-level block captures much finer speaker-related characteristics but less speaker-discriminative information. This phenomenon is consistent with a previous zero-shot study [28]. Please note that this visualization only analyzes the representation from global perspective, but the speaker representations used in the MTCR-VC are time-varying features.
Therefore, to further investigate the impact of different TCR blocks, we objectively evaluate the performance of MTCR-VC by sequentially discarding blocks from the high-level layer to the low-level layer. Specifically, the variant model w/o $Z_{s_{3}}$ only uses the first two TCR blocks in the speaker module and is trained from scratch to verify the role of $Z_{s_{3}}$, compared to the proposed MTCR-VC. Besides, discarding both second and third TCR blocks is also verified, forming the model w/o $Z_{s_{2}}$ $Z_{s_{3}}$. The validation results are presented in Table V. As can be seen, discarding the use of the third block in the model w/o $Z_{s_{3}}$ leads to decreased speaker accuracy. It suggests that the third block also learns time-varying speaker representation, which contains fine-grained speaker information, even though the averaged $Z_{s_{3}}$ shows little discrimination among different speakers from the global perspective in Fig 8. Furthermore, removing the second block and third block in the model w/o $Z_{s_{2}}$ $Z_{s_{3}}$ leads to more performance degradation in speaker accuracy. This indicates that the speaker representations learned in $Z_{s_{2}}$ and $Z_{s_{3}}$ are important for speaker accuracy in MTCR-VC. Additionally, it can be found that discarding the use of TCR blocks increases the WER. This outcome occurs because the speaker representation participates in the speech representation fusion, and the performance of speaker modeling affects the speech generation process.
(a) Spectrogram of a speaker utterance
(b) Pitch contour of the utterance
(c) Channel retrieval
(d) Temporal retrieval